OpenStudy (anonymous):

Write the following proportions as equivalent fractions. Do not solve A tree is 10 m tall has a shadow that is 50 m. What is the length of the shadow for a tree that is 5 m tall?

OpenStudy (anonymous):

\[\frac{ height }{ shadow }=\frac{ 10 }{ 50 }=\frac{ 5 }{ x }\]